# Training Large Language Models to Reason In Parallel with Global Forking Tokens

### v1

@@ 3.2 Evaluating SSFT on Reasoning Benchmarks
Pass@1 of Native Cons@6: Average performance of majority voting with 6 parallel generations
@@ 3.3 Evaluating Parallel Reasoning Diversity and Learned Matchings
Addressing RQ4, we show that our global forking tokens genuinely initiate distinct reasoning traces and offer a new mechanism for leveraging test-time compute.
Emerging Diverse Reasoning Modes. Using the $\mathrm{Cons@}6$ results in Table 1, we form six sets of generations, each prompted by a distinct ${\textnormal{g}}^{(i)}$. For each set, we show the average accuracy and the distribution of thinking-token counts: Figure 5 for SSFT with optimal bipartite matching and Figure 6 for SSFT with random matching. Reasoning length: Length partially indicates the diversity in reasoning, and we see clear differences for SSFT with optimal matching, despite the absence of hand-crafted matching rule or information about these traces. The consistency of these distributions across AIME24 and AIME25 indicates the differences is not from randomness, whereas randomly assigning a <thinki>, as in concurrent work (Wen et al., 2025), does not yield clear or consistent differences in reasoning length, as shown in Figure 6. Performance: After finetuning with random matching, prompting with a distinct <thinki> shows no meaningful impact ($\approx 61\%$ on AIME24 and $\approx 55\%$ on AIME25). With optimal matching, SSFT elicits distinct reasoning modes initiated by $\texttt{<think1>},\dots,\texttt{<think4>}$ that reach around $65\%$ on AIME24 and $\geq 59\%$ on AIME25, with different lengths. Although <think5> and <think6> are weaker due to shorter reasoning modes, the average of these and, especially, $\mathrm{Cons@}6$ performance improve consistently with them.
@@
For interpretability, we fine-tune using the same four teacher models for each $\mathbf{x}$}: GPT-OSS high, medium, R1, and Gemini. We call this dataset s1k-4teachers-reasoning dataset, and ask whether SSFT associate a unique ${\textnormal{g}}^{(i)}$ to the Gemini and R1 traces which have easily identifiable reasoning patterns, and then have the rest matched to OSS-high/OSS-medium traces in different ways (i.e. only matched to OSS-high, only matched to OSS-med, and matched to both). We track source indices only for evaluation; the model still receives an unordered set of traces with no source information for each question. We study three bipartite matching settings with SSFT. Figure 7 shows the evolution of matchings learned under these three hyperparameters: (a) random matching with four reserved ${\textnormal{g}}^{(i)}$, (b) optimal matching with four reserved ${\textnormal{g}}^{(i)}$, and (c) optimal matching with six reserved ${\textnormal{g}}^{(i)}$. Initially, all the configurations $\bm{\sigma}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ accumulate mass as there is no correlations between $\bm{g}$} sand $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$. Figure 7(a) shows SSFT under random matching does not shrink the size of configurations computed as optimal, meaning that no correlations are learned between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. By contrast, Figure 7(b) and Figure 7(c) show the emergence of only a strict subset of matching configurations in $\bm{\mathfrak{S}}_{\mathrm{P}}$. This indicates some correlations between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ are indeed learned through SSFT. We first visualize the learned matchings of the model with four ${\textnormal{g}}^{(i)}$ in Figure 8. We observe that ${\textnormal{g}}^{(1)}$ and ${\textnormal{g}}^{(2)}$ are uniquely matched to the R1 and Gemini traces, showing that SSFT can indeed uniquely associate ${\textnormal{g}}^{(i)}$ to sufficiently diverse reasoning traces. Now to confirm our previous hypothesis, we see the unique learned matchings $({\textnormal{g}}^{(3)},(\text{OSS-high, OSS-med})),({\textnormal{g}}^{(4)},\text{OSS-med})$. Furthermore, by connecting all the edges in $\bm{\mathfrak{S}}_{p}$ from SSFT with 6 forking tokens (Figure 7(c)), we also see unique learned matchings $({\textnormal{g}}^{(3)},\text{OSS-high}),({\textnormal{g}}^{(4)},\text{OSS-med}),({\textnormal{g}}^{(5)},(\text{OSS-med},\text{OSS-high}))$ in Figure 8. This confirms that the global forking tokens can identify unique correlations even among highly similar traces.
@@ 3.4 Ablation Study: Removing High Quality Small Dataset
To test whether our empirical gains are conditioned on the traces generated by our procedure and on a highly optimized small dataset (Muennighoff et al., 2025), we fine-tune on a public dataset that already provides sufficient reasoning traces per question: the 93k math set of Face (2025) (2-4 traces per question). Because this dataset has been successful for fine-tuning Qwen2.5-Math-7B, we adopt that base model and compare SSFT against SFT trained on all available traces. Details on the hyperparameters are in Appendix A.3.4. Addressing RQ5, Table 2 shows consistent improvements in both $\mathrm{Pass@}1$ and $\mathrm{Cons@}32$. The results indicate the SSFT is effective for larger sized public dataset with less diverse reasoning traces.
@@
Table 2: Performance of SSFT versus SFT trained solely on publicly available distillation targets. The setup uses the 93k math questions from Face (2025) with Qwen2.5-Math-7B as the base model. SFT-OpenR1-93k-7B uses the same distillation targets as SFT-mixed-distill-32B in Table 1.
@@ 4 Related work
Test-time Scaling. There has been a surge of work fine-tuning LLMs to reason longer, using reinforcement learning for frontier models (OpenAI, 2024; Shao et al., 2024; xAI, 2025; Yang et al., 2025a) and supervised fine-tuning for smaller ones (Muennighoff et al., 2025; Hu et al., 2025). These methods enable LLMs to improve reasoning by allocating more test-time compute to sequential, iterative refinement such as self-reflection (Guo et al., 2025; Liu et al., 2025). However, extended sequential reasoning can be more sensitive to the order of reasoning steps and may result in failures (Chen et al., 2024b), and performance can start to degrade beyond a certain length due to “overthinking” (Ghosal et al., 2025). Our goal is to study the effective use of diverse reasoning traces to fine-tune small language models, essential for agentic AI (Belcak et al., 2025).
@@
Parallel Reasoning. Parallel scaling methods such as self-consistency  (Wang et al., 2022) and Best-of-N (Lightman et al., 2023) improve LLM performance by generating multiple reasoning paths in parallel and aggregating them. These methods fundamentally require choosing a temperature that can generate diverse reasoning paths, but a recent theoretical work shows that increasing temperature can sometimes fail to increase diversity if language models are not trained towards coverage (Verine et al., 2025). Other search-based methods such as Monte Carlo tree search (MCTS) (Zhang et al., 2024) and Tree of Thoughts (ToT) (Yao et al., 2023) apply heuristic-guided search with an external verifier to do more deliberate search to increase the coverage. However, their dependence on heuristics and domain-specific knowledge can limit their applicable tasks. Regarding training LLMs with parallel reasoning traces,  Yang et al. (2025b) proposes training with parallel CoTs decomposed from sequential CoTs, and our concurrent work Wen et al. (2025) proposes to train with multiple reasoning traces distilled from teacher models. These works show native parallel scaling can surpass sequential scaling within certain token limits. Although we also show improvements in reasoning performance by simply fine-tuning with parallel reasoning traces, we empirically highlighted the limitation of manually assigning parallel think tags to target reasoning traces. If human assignment or random enforcement routes mismatched traces to a tag with a unique reasoning mode learned, the tag’s semantics blur due to bias in the ordering of parallel reasoning traces. This prevents tags from initiating distinct reasoning modes. We are the only work that leverages parallel think tags to construct a set-prediction loss via bipartite-matching optimization between tags and reasoning traces, turning the tags into global forking tokens. Empirically, we also improve $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ over baselines fine-tuned with parallel reasoning traces on the same dataset, whether on subsets or the full set. Another recent work (Zheng et al., 2025) encourages the use of parallel thinking tags by slightly upweighting rewards for correct trajectories that employ such tags during reinforcement learning. However, the cold-start SFT that instills parallel thinking from parallel reasoning traces is still performed without a set-prediction loss; as a result, the tags may initiate similar reasoning traces.
@@
Set-based Global Loss in Deep Learning. DETR introduces end-to-end object detection with a set global loss (Carion et al., 2020; Minderer et al., 2022), whose success in parallel bounding-box prediction inspires our approach. We are the first to extend this to language modeling: While DETR predicts a set of single tokens in parallel to match a list of bounding boxes, we predict a set of sequential thinking tokens, initiated by global forking tokens, and evaluate the matchings using autoregressive losses. We adapt set-based loss with autoregressive (AR) models, rather than diffusion-based models (Zhao et al., 2025; Nie et al., 2025; Ho et al., 2020), because AR models achieve superior reasoning performance. Several innovations are needed to make this loss practical for parallel reasoning with AR models. For example, to populate the cost matrix we need $\mathrm{N}\times\mathrm{M}$ forward passes, one for each combination of $\mathrm{N}$ think tags and $\mathrm{M}$ reasoning traces, whereas DETR requires only $\mathrm{N}$ forward passes for its $\mathrm{N}$ queries, since the same bounding-box predictions can be reused to compute matching costs against different ground-truth boxes. We therefore emphasize efficient training implementation in our extension, increasing training time by only $6.6\%$ compared to standard SFT (Appendix A.1).
@@ 5 Conclusion
Reproducibility. To ensure reproducibility of our work, we provide detailed hyper-parameter settings in A.3.2, and SSFT algorithm implementations in A.1. We are preparing the upcoming open-source release of our code and model.
@@ A.1 Algorithm: SSFT Implementation
13:   Compute optimal matching $\hat{\bm{\sigma}}_{k}$ between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}_{k}^{(j)}\}_{j=1}^{\mathrm{M}}$. Hungarian algorithm (Kuhn, 1955) can be applied to $\mathrm{C}$ to efficiently compute Equation 9 (Equation 3).
@@ A.3.1 Training Datasets
32B experiments with questions from s1(Main): We explain the process of generating our training dataset for experiments in Table 1, Figure 4, Figure 5, Figure 6, Figure 7. First, we use the 1000 questions from s1 (Muennighoff et al., 2025) and populate a pool of reasoning traces by distilling from GPT-OSS-120B-high reasoning, GPT-OSS-120B-medium reasoning, DeepSeek R1, Gemini Flash2.0 Thinking, and Claude4/4.1. We use temperature 1.0, maximum length of 32768, and sets high reasoning effort unless specified. We generate two traces per teacher model. We use Claude3.5 Sonnet to extract the answer from the distilled solutions and compare with the ground-truth answer. The correctness of these distilled traces are shown in 3.
@@ A.3.2 Training Hyperparameters
For consistency, we use Qwen2.5-32B-Instruct (Yang et al., 2025a) as the base model for all of our 32B experiments. We use standard fine-tuning hyperparameters: we train for 6 epochs with a global batch size of 32, which is derived from 4 gradient accumulation steps and distributed training with 8 GPUS ($4\times 8=32$). This results in 756 gradient steps. The maximum sequence length is set to 32,768. We train with bfloat16 with a learning rate of 1e51e-5 warmed up linearly for 5% and then decayed to 0 using a cosine schedule. We choose AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay 1e41e-4. We only backpropagates the completion loss, which is the loss on reasoning traces and the answers. Fine-tuning SSFT-32B plus loggings took 6.5 hours on 8 NVIDIA B200 GPUs using PyTorch FSDP, Liger Kernel (Hsu et al., 2024) for fused cross entropy loss, and FlashAttention-2 (Dao, 2023) for fused attention computation. Fine-tuning SSFT-32B (random $\bm{\sigma}$) took 6.3 hours, and Fine-tuning SFT-mixed-distill-32B took 6.1 hours. Even our baseline SFT-OSS-distill-32B with only one trace per question, and our attempt to reproduce s1.1 took 1.66 hours, which is longer than the time reported by Muennighoff et al. (2025). This is due to using 8 GPUs instead of 16 GPUs, hardware and package differences. When training with s1k-4mixed-reasoning, we added one extra epoch from 5 epochs to 6 epochs, since we have 4x reasoning traces, but we did not linearly increase the number of epochs, as these traces can be similar, and the number of distinct questions is still 1,000. Overall, we made sure all of our models are fine-tuned with consistent hyperapameters.

### v2

@@ 3.2 Evaluating SSFT on Reasoning Benchmarks
Average of Native Cons@6: Average performance of majority voting with 6 parallel generations
@@ 3.3 Evaluating Parallel Reasoning Diversity and Learned Matchings
Addressing RQ4, we show that our global forking tokens genuinely initiate distinct reasoning traces and offer a new mechanism for leveraging test-time compute. We call these tokens emergent because we do not manually identify distinct reasoning modes over the traces for different questions. For example, <think1> could lead to longer reasoning than <think5> for one question, but this may not be consistent for other similar questions. We do not add any auxiliary objective to skew the original language modeling objective.
Emerging Diverse Reasoning Modes. Using the $\mathrm{Cons@}6$ results in Table 1, we form six sets of generations, each prompted by a distinct ${\textnormal{g}}^{(i)}$. For each set, we show the average accuracy and the distribution of thinking-token counts: Figure 5 for SSFT with optimal bipartite matching and Figure 6 for SSFT with random matching. Reasoning length: Length partially indicates the diversity in reasoning, and we see clear differences for SSFT with optimal matching, despite the absence of hand-crafted matching rule or information about these traces. The consistency of these distributions across AIME24 and AIME25 indicates the differences is not from randomness, whereas randomly assigning a <thinki>, as in concurrent work (wen2025parathinker), does not yield clear or consistent differences in reasoning length, as shown in Figure 6. Performance: After finetuning with random matching, prompting with a distinct <thinki> shows no meaningful impact ($\approx 61\%$ on AIME24 and $\approx 55\%$ on AIME25). With optimal matching, SSFT elicits distinct reasoning modes initiated by $\texttt{<think1>},\dots,\texttt{<think4>}$ that reach around $65\%$ on AIME24 and $\geq 59\%$ on AIME25, with different lengths. Although <think5> and <think6> are weaker due to shorter reasoning modes, the average of these and, especially, $\mathrm{Cons@}6$ performance improve consistently with them.
@@
For interpretability, we fine-tune using the same four teacher models for each $\mathbf{x}$}: GPT-OSS high, medium, R1, and Gemini. We call this dataset s1k-4teachers-reasoning dataset, and ask whether SSFT associates a unique ${\textnormal{g}}^{(i)}$ to the Gemini and R1 traces which have easily identifiable reasoning patterns, and then have the rest matched to OSS-high/OSS-medium traces in different ways (i.e. only matched to OSS-high, only matched to OSS-med, and matched to both). We track source indices only for evaluation; the model still receives an unordered set of traces with no source information for each question. We study three bipartite matching settings with SSFT. Figure 7 shows the evolution of matchings learned under these three hyperparameters: (a) random matching with four reserved ${\textnormal{g}}^{(i)}$, (b) optimal matching with four reserved ${\textnormal{g}}^{(i)}$, and (c) optimal matching with six reserved ${\textnormal{g}}^{(i)}$. Initially, all the configurations $\bm{\sigma}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ accumulate mass as there is no correlations between $\bm{g}$} sand $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$. Figure 7(a) shows SSFT under random matching does not shrink the size of configurations computed as optimal, meaning that no correlations are learned between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. By contrast, Figure 7(b) and Figure 7(c) show the emergence of only a strict subset of matching configurations in $\bm{\mathfrak{S}}_{\mathrm{P}}$. This indicates some correlations between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ are indeed learned through SSFT. We first visualize the learned matchings of the model with four ${\textnormal{g}}^{(i)}$ in Figure 8. We observe that ${\textnormal{g}}^{(1)}$ and ${\textnormal{g}}^{(2)}$ are uniquely matched to the R1 and Gemini traces, showing that SSFT can indeed uniquely associate ${\textnormal{g}}^{(i)}$ to sufficiently diverse reasoning traces. Now to confirm our previous hypothesis, we see the unique learned matchings $({\textnormal{g}}^{(3)},(\text{OSS-high, OSS-med})),({\textnormal{g}}^{(4)},\text{OSS-med})$. Furthermore, by connecting all the edges in $\bm{\mathfrak{S}}_{p}$ from SSFT with 6 forking tokens (Figure 7(c)), we also see unique learned matchings $({\textnormal{g}}^{(3)},\text{OSS-high}),({\textnormal{g}}^{(4)},\text{OSS-med}),({\textnormal{g}}^{(5)},(\text{OSS-med},\text{OSS-high}))$ in Figure 8. This confirms that the global forking tokens can identify unique correlations even among highly similar traces.
@@ 3.4 Ablation Study: Removing High Quality Small Dataset
To test whether our empirical gains are conditioned on the traces generated by our procedure and on a highly optimized small dataset (muennighoff2025s1_simpletesttime), we fine-tune on a public dataset that already provides sufficient reasoning traces per question: the 93k math set of openr1 (2-4 traces per question). Because this dataset has been successful for fine-tuning Qwen2.5-Math-7B, we adopt that base model and compare SSFT against SFT trained on all available traces. Details on the hyperparameters are in Appendix A.3.4. Addressing RQ5, Table 2 shows consistent improvements in both $\mathrm{Pass@}1$ and $\mathrm{Cons@}32$. The results indicate the SSFT is effective for larger sized public dataset with less diverse reasoning traces.
@@
Table 2: Performance of SSFT versus SFT trained solely on publicly available distillation targets. The setup uses the 93k math questions from openr1 with Qwen2.5-Math-7B as the base model. SFT-OpenR1-93k-7B uses the same distillation targets as SFT-mixed-distill-32B in Table 1.
@@ 4 Related work
Test-time Scaling. There has been a surge of work fine-tuning LLMs to reason longer, using reinforcement learning for frontier models (o1; shao2024deepseekmathpushinglimitsmathematical; xAI2025; yang2025qwen3) and supervised fine-tuning for smaller ones (muennighoff2025s1_simpletesttime; hu2025distillationwhydistillation). These methods enable LLMs to improve reasoning by allocating more test-time compute to sequential, iterative refinement such as self-reflection (guo2025deepseekR1; liu2025understanding_r1zero). However, extended sequential reasoning can be more sensitive to the order of reasoning steps and may result in failures (chen2024premiseOrderMatters), and performance can start to degrade beyond a certain length due to “overthinking” (ghosal2025doesoverthinkinghelp). Our goal is to study the effective use of diverse reasoning traces to fine-tune small language models, essential for agentic AI (belcak2025small). Related to our work on aggregating diverse reasoning modes, yu2025chainofreasoning_towardsunifedmathematical; zheng2025learningtoreasonviaMixtureOfThought proposed incorporating three specific reasoning modes: natural language, code-based, and symbolic reasoning.
@@
Parallel Reasoning. Parallel scaling methods such as self-consistency  (wang2022selfconsistency) and Best-of-N (best_of_n) improve LLM performance by generating multiple reasoning paths in parallel and aggregating them. These methods fundamentally require choosing a temperature that can generate diverse reasoning paths, but a recent theoretical work shows that increasing temperature can sometimes fail to increase diversity if language models are not trained towards coverage (verine2025improvingwhentemperaturefailschangetheloss). Other search-based methods such as Monte Carlo tree search (MCTS) (zhang2024accessingmontecarlo) and Tree of Thoughts (ToT) (yao2023treeofthoughts) apply heuristic-guided search with an external verifier to do more deliberate search to increase the coverage. However, their dependence on heuristics and domain-specific knowledge can limit their applicable tasks. Regarding training LLMs with parallel reasoning traces,  yang2025multiverse proposes training with parallel CoTs decomposed from sequential CoTs, and our concurrent work wen2025parathinker proposes to train with multiple reasoning traces distilled from teacher models. These works show native parallel scaling can surpass sequential scaling within certain token limits. Although we also show improvements in reasoning performance by simply fine-tuning with parallel reasoning traces, we empirically highlighted the limitation of manually assigning parallel think tags to target reasoning traces. If human assignment or random enforcement routes mismatched traces to a tag with a unique reasoning mode learned, the tag’s semantics blur due to bias in the ordering of parallel reasoning traces. This prevents tags from initiating distinct reasoning modes. We are the only work that leverages parallel think tags to construct a set-prediction loss via bipartite-matching optimization between tags and reasoning traces, turning the tags into global forking tokens. Empirically, we also improve $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ over baselines fine-tuned with parallel reasoning traces on the same dataset, whether on subsets or the full set. Another recent work (zheng2025parallelr1towardsparallelthinkingviaRL) encourages the use of parallel thinking tags by slightly upweighting rewards for correct trajectories that employ such tags during reinforcement learning. However, the cold-start SFT that learns parallel thinking from diverse reasoning traces does not incorporate a set-prediction loss, which is necessary to prevent collapse in reasoning diversity.
@@
Set-based Global Loss in Deep Learning. DETR introduces end-to-end object detection with a set global loss (carion2020end_detr; owl_vit), whose success in parallel bounding-box prediction inspires our approach. We are the first to extend this to language modeling: While DETR predicts a set of single tokens in parallel to match a list of bounding boxes, we predict a set of sequential thinking tokens, initiated by global forking tokens, and evaluate the matchings using autoregressive losses. We adapt set-based loss with autoregressive (AR) models, rather than diffusion-based models (zhao2025d1; nie2025largediffusionmodels; DDPM), because AR models achieve superior reasoning performance. Several innovations are needed to make this loss practical for parallel reasoning with AR models. For example, to populate the cost matrix we need $\mathrm{N}\times\mathrm{M}$ forward passes, one for each combination of $\mathrm{N}$ think tags and $\mathrm{M}$ reasoning traces, whereas DETR requires only $\mathrm{N}$ forward passes for its $\mathrm{N}$ queries, since the same bounding-box predictions can be reused to compute matching costs against different ground-truth boxes. We therefore emphasize efficient training implementation in our extension, increasing training time by only $6.6\%$ compared to standard SFT (Appendix A.1).
@@ 5 Conclusion
Acknowledgment
This research was supported by compute resources provided by Amazon Web Services.
@@ A.1 Algorithm: SSFT Implementation
13:   Compute optimal matching $\hat{\bm{\sigma}}_{k}$ between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}_{k}^{(j)}\}_{j=1}^{\mathrm{M}}$. Hungarian algorithm (hungarian) can be applied to $\mathrm{C}$ to efficiently compute Equation 9 (Equation 3).
@@ A.3.1 Training Datasets
32B experiments with questions from s1(Main): We explain the process of generating our training dataset for experiments in Table 1, Figure 4, Figure 5, Figure 6, Figure 7. First, we use the 1000 questions from s1 (muennighoff2025s1_simpletesttime) and populate a pool of reasoning traces by distilling from GPT-OSS-120B-high reasoning, GPT-OSS-120B-medium reasoning, DeepSeek R1, Gemini Flash2.0 Thinking, and Claude4/4.1. We use temperature 1.0, maximum length of 32768, and sets high reasoning effort unless specified. We generate two traces per teacher model. We use Claude3.5 Sonnet to extract the answer from the distilled solutions and compare with the ground-truth answer. The correctness of these distilled traces are shown in 3.
@@ A.3.2 Training Hyperparameters
For consistency, we use Qwen2.5-32B-Instruct (yang2025qwen3) as the base model for all of our 32B experiments. We use standard fine-tuning hyperparameters: we train for 6 epochs with a global batch size of 32, which is derived from 4 gradient accumulation steps and distributed training with 8 GPUS ($4\times 8=32$). This results in 756 gradient steps. The maximum sequence length is set to 32,768. We train with bfloat16 with a learning rate of 1e51e-5 warmed up linearly for 5% and then decayed to 0 using a cosine schedule. We choose AdamW optimizer (loshchilov2017decoupled) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay 1e41e-4. We only backpropagates the completion loss, which is the loss on reasoning traces and the answers. Fine-tuning SSFT-32B plus loggings took 6.5 hours on 8 NVIDIA B200 GPUs using PyTorch FSDP, Liger Kernel (hsu2024ligerkernel) for fused cross entropy loss, and FlashAttention-2 (dao2023flashattention) for fused attention computation. Fine-tuning SSFT-32B (random $\bm{\sigma}$) took 6.3 hours, and Fine-tuning SFT-mixed-distill-32B took 6.1 hours. Even our baseline SFT-OSS-distill-32B with only one trace per question, and our attempt to reproduce s1.1 took 1.66 hours, which is longer than the time reported by muennighoff2025s1_simpletesttime. This is due to using 8 GPUs instead of 16 GPUs, hardware and package differences. When training with s1k-4mixed-reasoning, we added one extra epoch from 5 epochs to 6 epochs, since we have 4x reasoning traces, but we did not linearly increase the number of epochs, as these traces can be similar, and the number of distinct questions is still 1,000. Overall, we made sure all of our models are fine-tuned with consistent hyperapameters.